Calculate the mass of lead in 139g of Pb(NO3)2
1
Expert's answer
2020-11-02T09:05:19-0500

According to the chemical formula of Pb(NO3)2, where 1 formula unit contains 1 atom of lead, the number of the moles of lead nitrate and the number of the moles of lead are equal:

"n(Pb(NO_3)_2) =n(Pb)" .

The number of the moles of lead nitrate contained in 139 g is the ratio of its mass and its molar mass "M=331.21" g/mol:

"n(Pb(NO_3)_2) = \frac{m}{M} = \frac{139 \text{g}}{331.21\text{g/mol}} = 0.4197" mol.

Therefore, the number of the moles of lead is :

"n(Pb) = 0.4197" mol.

The molar mass of lead is 207.20 g/mol. Thus, its mass contained in 139 g of lead nitrate is:

"m(Pb) = nM = 0.4197\text{mol}\cdot207.20\text{g/mol}=86.96" , or 87 g.

Answer: 87 g of lead is contained in 139 g of Pb(NO3)2.